# Soft Phonon Charge-Density Wave Formation in the Kagome Metal KV3Sb5

### v1

@@ Abstract
Our findings raise the question as to why phonon softening is not observed in (Cs,Rb)V3Sb5 Li et al. (2021); Subires et al. (2023), for which there are several potential contributing factors. As the CDW transitions in (Cs,Rb)V3Sb5 are first-order Song et al. (2022); Zhang et al. (2024), the transition may occur before the corresponding phonons significantly soften, which contrasts with KV3Sb5 that shows a second-order CDW transition Scagnoli et al. (2024). Compared to KV3Sb5 that shows a single CDW phase in its temperature-pressure phase diagram Du et al. (2021); Kautzsch et al. (2023), CsV3Sb5 has CDWs modulated by $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},\frac{1}{4})$ at ambient pressure Hu et al. (2022b); Stahl et al. (2022); Xiao et al. (2023), and a pressure-induced $(0,\frac{3}{8},\frac{1}{2})$ CDW under pressure Zheng et al. (2022); Stier et al. (2024). In RbV3Sb5, there may also be a pressure-induced CDW distinct from the ambient pressure CDW, although its transport signatures are much weaker compared to CsV3Sb5 Wang et al. (2021); Du et al. (2022a). Such a coexistence of distinct CDW instabilities close in energy may frustrate the system and hinder the softening of phonons at a particular wave vector Chen et al. (2024). Nonetheless, as the CDW-related diffuse scattering in (Cs,Rb)V3Sb5 Subires et al. (2023) are similar to that in KV3Sb5, the presence of soft phonons in these compounds cannot be excluded. As this study on KV3Sb5 and previous IXS studies of (Cs,Rb)V3Sb5 are performed in different Brillouin zones, a comprehensive search of soft phonons across multiple zones in (Cs,Rb)V3Sb5 will be helpful to elucidate whether phonon softening is common to the CDW formation of AAV3Sb5.
@@
Given the role of EPC in driving the CDW of AAV3Sb5, it may also play a role in the superconductivity that emerges from the CDW state You et al. (2025). In such a scenario, the pairing symmetry is expected to be nodeless, for which there has been considerable evidence Duan et al. (2021); Zhou et al. (2024); Mine et al. (2024); Kaczmarek et al. (2024), although there are also reports of unconventional pairing Guguchia et al. (2023); Deng et al. (2024); Zhao et al. (2024). The role of the lattice (and thus EPC) in the superconductivity of AAV3Sb5 is also underscored by the tuning of $T_{\rm c}$ via structural crossovers or transitions under hydrostatic pressure Tsirlin et al. (2022); Yu et al. (2022); Du et al. (2022b).
@@
In conclusion, we find that the CDW in KV3Sb5 forms via phonon softening towards zero energy around $T_{\rm CDW}$, with soft phonons occurring over an extended momentum range along ALA-L but a narrow range along HLH-L. Such an in-plane anisotropy of soft phonons is reproduced in first-principles calculations, which arise from a momentum-dependent EPC. These results evidence a conventional mechanism of the CDW formation in KV3Sb5, similar to that in transition metal dichalcogenides, which likely also applies to other members of the AAV3Sb5 family.

### v2

@@ Abstract
Our findings raise the question as to why phonon softening is not observed in (Cs,Rb)V3Sb5 Li et al. (2021); Subires et al. (2023), for which there are several potential contributing factors. As the CDW transitions in (Cs,Rb)V3Sb5 are first-order Song et al. (2022); Zhang et al. (2024), the transition may occur before the corresponding phonons significantly soften, which contrasts with KV3Sb5 that shows a second-order CDW transition Scagnoli et al. (2024). Compared to KV3Sb5 that shows a single CDW phase in its temperature-pressure phase diagram Du et al. (2021); Kautzsch et al. (2023), CsV3Sb5 has CDWs modulated by $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},\frac{1}{4})$ at ambient pressure Hu et al. (2022a); Stahl et al. (2022); Xiao et al. (2023), and a pressure-induced $(0,\frac{3}{8},\frac{1}{2})$ CDW under pressure Zheng et al. (2022); Stier et al. (2024). In RbV3Sb5, there may also be a pressure-induced CDW distinct from the ambient pressure CDW, although its transport signatures are much weaker compared to CsV3Sb5 Wang et al. (2021); Du et al. (2022b). Such a coexistence of distinct CDW instabilities close in energy may frustrate the system and hinder the softening of phonons at a particular wave vector Chen et al. (2025). Nonetheless, as the CDW-related diffuse scattering in (Cs,Rb)V3Sb5 Subires et al. (2023) are similar to that in KV3Sb5, the presence of soft phonons in these compounds cannot be excluded. As the soft phonons may be weak in some Brillouin zones [e.g. soft phonons in KV3Sb5 were not detected at ${\bf Q}=(3.5,0,0.5)$], a comprehensive search of soft phonons across multiple zones in (Cs,Rb)V3Sb5 will be helpful to elucidate whether phonon softening is common to the CDW formation of AAV3Sb5.
@@
Given the role of EPC in driving the CDW of AAV3Sb5, it may also play a role in the superconductivity that emerges from the CDW state Jiang et al. (2023); You et al. (2025). In such a scenario, the pairing symmetry is expected to be nodeless, for which there has been considerable evidence Duan et al. (2021); Zhou et al. (2024); Mine et al. (2024); Kaczmarek et al. (2025), although there are also reports of unconventional pairing Guguchia et al. (2023); Deng et al. (2024); Zhao et al. (2024). The role of the lattice (and thus EPC) in the superconductivity of AAV3Sb5 is also underscored by the tuning of $T_{\rm c}$ via structural crossovers or transitions under hydrostatic pressure Tsirlin et al. (2022); Yu et al. (2022); Du et al. (2022a). Furthermore, our calculations show that EPC reasonably accounts for superconductivity of KV3Sb5, when its CDW is destabilized under pressure 1.
@@
In conclusion, we find that the CDW in KV3Sb5 forms via phonon softening towards zero energy around $T_{\rm CDW}$, with soft phonons observed over an extended momentum range along ALA-L but a narrow range along HLH-L. The phonon softening of KV3Sb5 is reproduced in first-principles calculations, and is shown to arise from a momentum-dependent EPC. These results evidence a conventional EPC mechanism of the CDW formation in KV3Sb5, similar to that in transition metal dichalcogenides, which likely also applies to other members of the AAV3Sb5 family.
@@
Note: After the submission of this work, soft phonons were also reported for CsV3Sb5 McGuinness et al. (2025), pointing to a common soft-phonon CDW formation in the AAV3Sb5 compounds.